Session: Sep 1999 (session 3)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Torun (K.Borkowski at Tue Sep 21 11:23:58 METDST 1999) Minor failures
The observations stopped at 6:41 UT with the power failure. At the time of writing the telescope is still inoperative due to ongoing repair of a damage caused by the power failure.

Medicina (Daniele Dallacasa at Tue Sep 21 11:36:48 METDST 1999) Success
Observation completed without known problems.

Noto (C.Trigilio at Tue Sep 21 12:55:56 METDST 1999) Failure
Noto did not observe because of the failure of the recorder capstain motor

Westerbork (Tony Foley at Tue Sep 21 14:05:18 METDST 1999) Success
No known problems, but the online noise cal used is only very rough. The single dish setup was used (RT7)

Effelsberg (A. Kraus at Wed Sep 22 09:27:53 METDST 1999) Success

Onsala ( at Sun Sep 26 11:18:08 METDST 1999) Success

Jodrell (Alastair Gunn at Wed Sep 29 13:27:45 METDST 1999) Success
No known problems.

Calibration (Jean\-Francois Desmurs at Sun Nov 7 15:47:11 MET 1999) Success
The ANTAB formatted file for this experiment can be found on VLBEER in the sep99 directory under the name: v016b2.antab. Read the comments in the file to get further information. (Rem : Jb most of Tsys values are bad )